OpenStudy (anonymous):

In the box, complete the first 4 steps for graphing the quadratic function given. (Use ^ on the keyboard to indicate an exponent.) Then print a sheet of graph paper and graph the quadratic function to turn in to your teacher. Be sure to label the axes and vertex. y = -x 2 - 4x - 3 ALGEBRA 1

OpenStudy (anonymous):

@Andresfon12

OpenStudy (anonymous):

so far to find the x axis is (b)/2a

OpenStudy (anonymous):

\[\frac{ b }{ 2a }\]

OpenStudy (anonymous):

here is a site how to find the vertex http://www.purplemath.com/modules/sqrvertx.htm

OpenStudy (anonymous):

@Andresfon12 Is this it? Multiple sites give me different answers y= -x^2 - 4x - 3 Completing Square to put into Vertex Form y = -(x+2)^2 + 4 - 3 y = -(x+2)^2 + 1 ||Parabola opens downward (-1 <0) along x = -2 , V(-2,1)

OpenStudy (anonymous):

first thing we know that x axis b=-4 and a=-1 so -4/2(-1)

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20120604092133AAuMo9t This link has multiple answers but I'm just stumped on this. It's the last question on my homework and all the other questions weren't like this

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20120604092133AAuMo9t is right forget some of this lol

OpenStudy (anonymous):

The Best Answer by OldTeacher?